(a) The pitch circle diameter of the smaller of the two spur wheels which mesh externally an lute teeth is 100 mm. The number of teeth are 16 and 32. The pressure angle is 20 degrees a endum is 0.32 of the circular pitch. Find the length of the path of contact of the pair of teeth.


A pair of gears, having 40 and 30 teeth respectively are of 25° involute form. The addendum le m and the module pitch is 2.5 mm. If the smaller wheel is the driver and rotates at 1500 r.p.n velocity of sliding at the point of engagement and at the point of disengagement.


1
Expert's answer
2021-11-17T10:44:01-0500

length of the path of contact of the pair of teeth = length of arc of contact × "\cos\theta"


"\theta = 20°"

length of arc of contact = 2 × pc

pc = 2π ×100 = 200π mm


"\therefore" length of arc of contact = 2 × 200π = 400π


length of the path of contact of the teeth = 400π × cos20 = 1180.9mm "\approx" 1.18m"."